The following data are expected to follow a linear relation of the form y = ax+b. Obtain the best linear relation in accordance with a least-squares analysis. Calculate the standard deviation of the data from the predicted straight-line relation :

x| 0.9| 2.3| 3.3| 4.5| 5.7 |6.7

y| 1.1 |1.6| 2.6| 3.2| 4.0| 5.0
Answer #1

Following table shows the calcualtion for regression line:

X Y X^2 Y^2 XY
0.9 1.1 0.81 1.21 0.99
2.3 1.6 5.29 2.56 3.68
3.3 2.6 10.89 6.76 8.58
4.5 3.2 20.25 10.24 14.4
5.7 4 32.49 16 22.8
6.7 5 44.89 25 33.5
Total 23.4 17.5 114.62 61.77 83.95

Number of data points : n= 6

Now

S_{yy}=\sum y^{2}-\frac{\left (\sum y \right )^{2}}{n}=10.7283

S_{xx}=\sum x^{2}-\frac{\left (\sum x \right )^{2}}{n}=23.36

S_{xy}=\sum xy-\frac{\left (\sum x \right )\left (\sum y \right )}{n}=15.7

Slope of the regression equation is

a=b_{1}=\frac{S_{xy}}{S_{xx}}=0.6721

and intercept of the equation will be

b=b_{0}=\frac{1}{n}(\sum y - b_{1} \sum x)=0.2955

So least sqaure line between the data is

y = 0.6721 x + 0.2955

--------------

Let us find SSE first :

SSE=\sum y^{2}-b_{0}\sum y-b_{1}\sum xy=0.1765

So standard error of estimate will be

SE=\sqrt{\frac{SSE}{n-2}}=0.2101